# Quantum walks, quantum cellular automata, quantum field theory, symmetry, interactions

### v1

@@ 7.4.2 Toeplitz Matrices and Their Eigenvalues
We can check that $\tilde{f}$ is convergent and real, and has the properties $\tilde{f}(k)=\tilde{f}(-k)$, $\tilde{f}(k+2\pi)=\tilde{f}(k)$. Using the above observation connecting the matrix elements of $\bar{D}$ to the Taylor expansion of the arctangent, we can explicitly calculate
@@
which produces a striking square wave, as shown in Fig. 2.
Figure 2: The discrete Fourier transform $\tilde{f}(k)$ of the function $f(n)$ defining our family of Toeplitz matrices.
The bound on the smallest eigenvalue follows from the following theorems.
Theorem 1 Parter61 ; Serra98 . If $\{A_{M}\}$ is a sequence of $M\times M$ self-adjoint Toeplitz matrices (defined by a function $f(n)$ with well-defined transform $\tilde{f}(k)$), and if $\sigma_{M}$ is the minimum singular value of $A_{M}$, then $|\sigma_{M}-\sigma_{\infty}|<e^{-\alpha M}$ for some $\alpha>0$. (The smallest singular value of a self-adjoint matrix is the smallest absolute value of an eigenvalue, $|\lambda|$. Since the matrices $\{A_{M}\}$ are positive, this is the same as the smallest eigenvalue.)
Theorem 2 Parter61 ; Serra98 . The limit $\sigma_{\infty}$ of the smallest singular value of the matrix family $\{A_{M}\}$ is $\min_{k}\tilde{f}(k)$.
As we can see in Fig. 2, the minimum value of $\tilde{f}(k)$ is 0. So the minimum eigenvalues of the matrices $\bar{D}$ approach 0 exponentially quickly as $M$ becomes large. (Note that the form of $\tilde{f}(k)$ also implies that the maximum eigenvalue approaches 1 exponentially quickly from below.) While we do not have an analytical expression for $\alpha$, we can estimate it numerically.
@@ 7.4.3 Numerical Results
We have numerically calculated the minimum eigenvalue for the $(M+1)\times(M+1)$ matrix $\bar{D}$ for an interaction range up to $M=100$. As shown in Fig. 3, these eigenvalue drop exponentially with the range $M$. (Because these eigenvalues are so tiny, high-precision numerical calculations are required, with up to 110 digits of precision.) Numerical fitting suggests that these eigenvalues fall off as $e^{-\alpha M}$ for $\alpha\approx 1.755$. (We note that this is close to $\sqrt{\pi}$, but in the absence of an analytical formula this is likely to be a coincidence.)
Figure 3: The smallest eigenvalue of the $(M+1)\times(M+1)$ matrix $\bar{D}$ vs. the interaction range $M$. These values were calculated using high-precision numerics.
@@
The behavior of this eigenvalue tells us two things about locality and negative energies. First, the coupling to negative energy modes is nonzero for every finite range. Any pair of coupled QCAs of this sort will have an amplitude to produce negative-energy bosons; to eliminate this amplitude requires non-vanishing interactions at arbitrary distances. But second, since the minimum eigenvalue falls off so rapidly with range, it is possible to suppress these amplitudes to the point where they are essentially negligible over any desired length of time. If we let the lattice size $N$ become very large, with a very small lattice spacing $\Delta x$ we can have the interaction range $M\Delta x$ be imperceptibly small; effectively indistinguishable from a point interaction at longer length scales. If we take the limit $\Delta x\rightarrow 0$ and $N\rightarrow\infty$ to recover QFT, these become zero-range interactions.
Figure 4: The absolute value $|v_{j}|$ of the $j$th component of the eigenvector corresponding to the smallest eigenvalue of the $(M+1)\times(M+1)$ matrix $\bar{D}$ for $M=100$. Numerical fitting shows that it is extremely close to an exact Gaussian curve.
@@
In fact, this near-locality is even stronger than that. In Fig. 4 we plot the absolute value of the eigenvector component for the minimum value of $\bar{D}$ when $M=100$. To the eye it appears to be a near-perfect Gaussian. Numerical fitting shows that it is, in fact, extremely close to a perfect Gaussian. This means that the amplitude to produce negative-energy bosons can be exponentially suppressed by an interaction over a finite range that also falls off rapidly away from the center.
@@
One final note about the accuracy of these numerical calculations. The form of the matrix used to calculate the results shown in Figs. 3 and 4 is given by the large $N$ approximation in Eq. (108), and is exact only in the limit $N\rightarrow\infty$. However, any corrections should scale like $O((M/N)^{2})$, and so also be extremely small.
8 Conclusion
@@
We have shown in this paper that it is possible to construct a QCA that recovers free QED in the long-wavelength limit. In the case of bosons, this is accomplished by making the degrees of freedom of the local sites harmonic oscillator modes, and coupling them by alternating quadratic interactions. This construction produces the Maxwell QFT by using a six-dimensional internal space. The result is a two-dimensional positive-energy sector (representing the two helicity states), a two-dimensional negative-energy counterpart, and a two-dimensional zero-energy sector that is static and decoupled from the rest.
@@
Using oscillator modes makes the HIlbert space of the local sites infinite dimensional. This is an interesting comparison with our earlier derivation of QCAs with fermionic statistics in 2 or more spatial dimensions, which required us to confine the state to the totally antisymmetric subspace of a QCA with high local dimension. It seems that to recover fermionic or bosonic statistics requires high-dimensional subsystems, albeit for different reasons.
@@
The existence of negative-energy solutions for both the fermionic and bosonic cases follows from the time-reversal symmetry of the QCAs. This creates a dilemma, since a purely local interaction between two QCAs will then always be able to excite negative-energy bosons, which could produce a cascade of particle-antiparticle creation. We examined this problem in a 1D model and showed that amplitudes for negative-energy states cannot be eliminated by interactions with finite range, but they can be suppressed without limit, with interactions that appear effectively local at longer length scales.
@@
This demonstration that the amplitude of negative-energy solutions can be suppressed exponentially in the interaction range was shown for a particular form of interaction in a 1D model. Numerical studies are more challenging in higher spatial dimensions, since the size of the analogous matrix will grow like a higher power of the range. It also will no longer be a Toeplitz matrix, which may complicate the analysis. While we do not know, we expect that the suppression will most likely be even greater in higher dimensions, given the larger number of degrees of freedom available. It also raises the question of how zero-energy degrees of freedom, which arise in the bosonic QCA models derived above in higher dimensions, play a role in interacting theories. This is a subject for future research.
@@
TAB and LM contributed equally to this work.

### v2

@@ 7.4.2 Toeplitz Matrices and Their Eigenvalues
We can check that $\tilde{f}$ is convergent and real and has the properties $\tilde{f}(k)=\tilde{f}(-k)$ and $\tilde{f}(k+2\pi)=\tilde{f}(k)$. Using the above observation connecting the matrix elements of $\bar{D}$ to the Taylor expansion of the arctangent, we can explicitly calculate
@@
which produces a striking square wave, as shown in Figure 2.
Figure 2: The discrete Fourier transform $\tilde{f}(k)$ of the function $f(n)$ defining our family of Toeplitz matrices. (Note, these are dimensionless numbers.)
The bound on the smallest eigenvalue follows from the following theorems:
{Theorem}
[Parter61 ; Serra98 ] If $\{A_{M}\}$ is a sequence of $M\times M$ self-adjoint Toeplitz matrices (defined by a function $f(n)$ with well-defined transform $\tilde{f}(k)$), and if $\sigma_{M}$ is the minimum singular value of $A_{M}$, then $|\sigma_{M}-\sigma_{\infty}|<e^{-\alpha M}$ for some $\alpha>0$. (The smallest singular value of a self-adjoint matrix is the smallest absolute value of an eigenvalue, $|\lambda|$. Since the matrices $\{A_{M}\}$ are positive, this is the same as the smallest eigenvalue.)
{Theorem}
[Parter61 ; Serra98 ] The limit $\sigma_{\infty}$ of the smallest singular value of the matrix family $\{A_{M}\}$ is $\min_{k}\tilde{f}(k)$.
As we can see in Figure 2, the minimum value of $\tilde{f}(k)$ is 0. So the minimum eigenvalues of the matrices $\bar{D}$ approach 0 exponentially quickly as $M$ becomes large. (Note that the form of $\tilde{f}(k)$ also implies that the maximum eigenvalue approaches 1 exponentially quickly from below.) While we do not have an analytical expression for $\alpha$, we can estimate it numerically.
@@ 7.4.3 Numerical Results
We have numerically calculated the minimum eigenvalue for the $(M+1)\times(M+1)$ matrix $\bar{D}$ for an interaction range up to $M=100$. As shown in Figure 3, these eigenvalues drop exponentially with the range $M$. (Because these eigenvalues are so small, high-precision numerical calculations are required, with up to 110 digits of precision.) Numerical fitting suggests that these eigenvalues fall off as $e^{-\alpha M}$ for $\alpha\approx 1.755$. (We note that this is close to $\sqrt{\pi}$, but in the absence of an analytical formula, this is likely to be a coincidence.)
@@
The behavior of this eigenvalue tells us two things about locality and negative energies. First, the coupling to negative-energy modes is non-zero for every finite range. Any pair of coupled QCAs of this sort will have an amplitude to produce negative-energy bosons; to eliminate this amplitude requires non-vanishing interactions at arbitrary distances. But second, since the minimum eigenvalue falls off so rapidly with range, it is possible to suppress these amplitudes to the point where they are essentially negligible over any desired length of time. If we let the lattice size $N$ become very large, with a very small lattice spacing $\Delta x$, we can have the interaction range $M\Delta x$ be imperceptibly small—effectively indistinguishable from a point interaction at longer length scales. If we take the limit $\Delta x\rightarrow 0$ and $N\rightarrow\infty$ to recover QFT, these become zero-range interactions.
Figure 3: The smallest eigenvalue of the $(M+1)\times(M+1)$ matrix $\bar{D}$ vs. the interaction range $M$. These values were calculated using high-precision numerics. (Note, these are dimensionless numbers.)
@@
In fact, this near-locality is even stronger than that. In Figure 4, we plot the absolute value of the eigenvector component for the minimum value of $\bar{D}$ when $M=100$. To the eye, it appears to be a near-perfect Gaussian. Numerical fitting shows that it is, in fact, extremely close to a perfect Gaussian. This means that the amplitude to produce negative-energy bosons can be exponentially suppressed by an interaction over a finite range that also falls off rapidly away from the center.
Figure 4: The absolute value $|v_{j}|$ of the $j$th component of the eigenvector corresponding to the smallest eigenvalue of the $(M+1)\times(M+1)$ matrix $\bar{D}$ for $M=100$. Numerical fitting shows that it is extremely close to an exact Gaussian curve. (Note that, since this is an eigenvector, it can be multiplied by an arbitrary overall factor.)
@@
One final note about the accuracy of these numerical calculations. The form of the matrix used to calculate the results shown in Figures 3 and 4 is given by the large $N$ approximation in Equation (109) and is exact only in the limit $N\rightarrow\infty$. However, any corrections should scale like $O((M/N)^{2})$ and so also be extremely small.
8 Conclusions
The correspondence between quantum cellular automata (QCA) and quantum field theory (QFT) reveals a compelling perspective: that the fundamental laws of physics may emerge from an underlying information-processing algorithm. This framework—where QFTs arise from quantum computations on a discrete spacetime lattice—holds deep conceptual appeal. It provides a potential origin story for physical law itself. It offers a path toward addressing persistent challenges in continuum QFTs, such as renormalization, by embedding field dynamics in a fundamentally finite structure. It opens avenues for simulating QFTs on quantum hardware. And, most significantly, it raises the provocative possibility that the universe’s own fabric may be inherently discrete and computational.
@@ 8 Conclusions
We have shown in this paper that it is possible to construct a QCA that recovers free QED in the long-wavelength limit. In the case of bosons, this is accomplished by making the degrees of freedom of the local sites harmonic oscillator modes and coupling them by alternating quadratic interactions. This construction produces the Maxwell QFT by using a six-dimensional internal space. The result is a two-dimensional positive-energy sector (representing the two helicity states), a two-dimensional negative-energy counterpart, and a two-dimensional zero-energy sector that is static and decoupled from the rest.
@@
Using oscillator modes makes the Hilbert space of the local sites infinite-dimensional. This is an interesting comparison with our earlier derivation of QCAs with fermionic statistics in two or more spatial dimensions, which required us to confine the state to the totally antisymmetric subspace of a QCA with high local dimension. It seems that to recover fermionic or bosonic statistics requires high-dimensional subsystems, albeit for different reasons.
@@
To recover QFTs describing physical systems requires interactions. Such QCAs have been treated in a number of recent works Arrighi20 ; MlodinowBrun21 ; Eon23 ; Centofanti24 . One problem in adding interactions is the existence of negative-energy states. Negative-energy solutions for both the fermionic and bosonic cases follow from the time-reversal symmetry of the QCAs. This creates a dilemma, since a purely local interaction between two QCAs will then always be able to excite negative-energy bosons, which could produce a cascade of particle–anti-particle creation. We examined this problem in a 1D model and showed that amplitudes for negative-energy states cannot be eliminated by interactions with a finite range, but they can be suppressed without limit, with interactions that appear effectively local at longer length scales.
@@
This demonstration that the amplitude of negative-energy solutions can be suppressed exponentially in the interaction range was shown for a particular form of interaction in a 1D model. Numerical studies are more challenging in higher spatial dimensions, since the size of the analogous matrix will grow like a higher power of the range. It will also no longer be a Toeplitz matrix, which may complicate the analysis. While we do not know, we expect that the suppression will most likely be even greater in higher dimensions, given the larger number of degrees of freedom available. It also raises the question of how zero-energy degrees of freedom, which arise in the bosonic QCA models derived above in higher dimensions, play a role in interacting theories. Could there also be other mechanisms to suppress the coupling to negative-energy (and zero-energy) modes, besides those discussed in this paper? Or could these modes end up playing a nontrivial role in a 3D theory? These are subjects for future research.
@@
Conceptualization, T.A.B. and L.M.; software, T.A.B.; validation, T.A.B. and L.M.; formal analysis, T.A.B. and L.M.; resources, T.A.B. and L.M.; data curation, N/A; writing, T.A.B. and L.M. contributed equally to all aspects of the writing, review and editing; visualization, T.A.B.; supervision, N/A; project administration, T.A.B.; funding acquisition, T.A.B.. All authors have read and agreed to the published version of the manuscript.